OpenStudy (gabylovesyou):

What is the equation in point-slope form of the line passing through (−2, 1) and (4, 13)? y − 1 = −2(x − 4) y − 1 = 2(x + 2) y − 13 = −2(x + 2) y − 13 = 2(x + 4)

OpenStudy (anonymous):

Can you calculate the slope of the line containing these points using \[m=\frac{ y_2 - y_1 }{ x_2 - x_1 }\]

OpenStudy (gabylovesyou):

2

OpenStudy (anonymous):

Great. Now, in slope-intercept form you have\[y=2x+b\]Substitute the x and y coordinates from either one of the given points into this equation and solve for b. Done

OpenStudy (gabylovesyou):

what do u mean ?

OpenStudy (anonymous):

Choose either one of the given points, let's say (-2, 1). The x-coordinate is -2 and the y-coordinate is 1. Substitute these values into the slope-intercept form, giving\[y=2x+b\]\[1=2\left( -2 \right)+b\]Are you now able to solve for b?

OpenStudy (vanessaperezxo):

y = 2x + b is the slope-intercept form. You have to fill in variables. So what you need to do is substitute the variable for either one of the points like @ospreytriple said.

OpenStudy (gabylovesyou):

b = 5

OpenStudy (anonymous):

Terrific. Now you have the value of m and of b, which is everything you need. Now just substitute these values for m and b into \[y=mx+b\] and you have your solution. BTW, it doesn't matter which of the two points you choose in the previous step, you will grt thr same answer.

OpenStudy (vanessaperezxo):

Yes, for the problem 1 = 2(-2) + b . That would equal 5.

OpenStudy (gabylovesyou):

ok im lost again

OpenStudy (vanessaperezxo):

ok, which part don't you understand?

OpenStudy (gabylovesyou):

how do i get my answer after finding b?

OpenStudy (anonymous):

\[y=mx+b\]with m=2 and b=5 gives\[y=2x+5\]which is the desired equation in slope-intercept form.

OpenStudy (gabylovesyou):

its not in my answers ?

OpenStudy (solomonzelman):

you just need to plug it into a point slope form

OpenStudy (anonymous):

Sorry. If there are no errors in the question you typed in originally, this is the correct answer.

OpenStudy (solomonzelman):

\(\large\color{slate}{ \displaystyle y-y_1=m\left(x-x_1\right)}\)

OpenStudy (gabylovesyou):

so its B ?

OpenStudy (solomonzelman):

your slope is 2 (one of) your point(s) is (-2, 1) plug this information into \(\large\color{slate}{ \displaystyle y-y_1=m\left(x-x_1\right)}\) where \(\large\color{slate}{ \displaystyle \left(x_1,y_1\right)}\) is the point. and \(\large\color{slate}{ \displaystyle m}\) is the slope.

OpenStudy (solomonzelman):

yuop

OpenStudy (solomonzelman):

it is B. \(\large\color{slate}{ \displaystyle y-1=2\left(x+2\right)}\)

OpenStudy (vanessaperezxo):

simplify the equation to see if it works. 

OpenStudy (vanessaperezxo):

but i agree the answer is B.

OpenStudy (vanessaperezxo):

It passes by the points (-2, 1) and (4,13)